# One Step is Enough: Multi-Agent Reinforcement Learning based on One-Step Policy Optimization for Order Dispatch on Ride-Sharing Platforms

### v1

@@ Conclusion
In this paper, we propose two novel MARL methods for ride-sharing that do not rely on estimating V-values or Q-values. First, we successfully adapt GRPO to the ride-sharing context, achieving promising performance by eliminating the estimation errors associated with the V-value network in PPO. Building on the insights from GRPO, we further introduce the OSPO method, which enables policy optimization using only single-step rewards. To the best of our knowledge, OSPO is the simplest and most efficient reinforcement learning method for order dispatch on ride-sharing platforms to date. Extensive experiments on real-world ride-sharing data demonstrate that our methods consistently outperform previous approaches in most scenarios, while requiring only a simple MLP network and exhibiting low GPU utilization.
Appendix
Reward Function Design
The reward function is shaped following previous work (Hu, Feng, and Li 2025):
$\text{r}(s_{i},u_{i})$ =$=\begin{cases}\beta_{1}+\beta_{2}p^{in}_{i}-\beta_{3}p^{out}_{i}-\beta_{4}\chi_{i}-\beta_{5}\rho_{i},|u_{i}|=1\\ 0,|u_{i}|=0\end{cases}$ (13)
where $\beta_{1}$ to $\beta_{5}$ are non-negative weights representing the platform’s valuation of each term, $p^{in}_{i}$ and $p^{out}_{i}$ represent the income from passengers and the payout to drivers, respectively. The variables $\chi_{i}$ and $\rho_{i}$ represent the number of en-route orders that will exceed their scheduled time and the additional travel time of all en-route orders when the assigned order is added to the scheduled route of agent $i$, respectively. This reward function is designed to comprehensively consider the interests of the platform, drivers, and passengers, mimicking the operation of a real-world ride-sharing platform. It is important to emphasize that $p^{in}_{i}$ and $p^{out}_{i}$ are calculated based on the order distance and the additional travel distance for the agent, respectively. When calculating travel time, we will utilize the Traveling Salesman Problem (TSP) to optimize the agent’s route.
Algorithm
@@
The detailed algorithm of our GRPO/OSPO for ride sharing method is illustrated in Algorithm 1.
Algorithm 1 GRPO for Ride Sharing
@@
1:  Initialize experience buffer $\mathcal{D}\leftarrow\emptyset$
2:  for episode $=1$ to $M$ do
@@
10:        Store transition ($(\{s_{i,t}\},s^{o}_{t},U_{t},\{r_{i,t+1}\},\{s_{i,t+1}\},s^{o}_{t+1})$ in $\mathcal{D}$
@@
12:     Policy Optimization:
13:     for epoch $=1$ to $M$ do
14:        Sample batch $\mathcal{B}\sim\mathcal{D}$
15:        Compute group advantage $\hat{A}_{t}$
16:        Update policy parameters $\theta$ by maximizing Eq. 1
17:     end for
18:     if current policy $\pi_{\theta}$ outperforms $\pi_{\phi}$ then
19:        Update best checkpoint: $\phi\leftarrow\theta$
20:     end if
21:  end for
Related Work
@@
In this section, we provide a brief overview of current ride-sharing methods. Starting from the Independent Double Deep Q-Learning (IDDQN)-based DeepPool (Al-Abbasi, Ghosh, and Aggarwal 2019), independent DTDE methods have become a mainstream paradigm, as the large number of agents, along with the extensive state and action spaces, severely hinders the development of centralized methods. Subsequent works have primarily aimed to enhance algorithms by fostering better cooperation among agents. For example, Graph Neural Networks (GNNs) have been widely explored to effectively utilize neighboring information (Hu, Feng, and Li 2025). Li et al. (Li et al. 2019) also propose using mean field approaches to capture more useful global information.
@@
Additionally, some works attempt to transfer the successes of CTDE and CTCE to ride-sharing, where centralized critics can facilitate better cooperation among agents. However, these approaches face various shortcomings and challenges. Lima et al. (De Lima et al. 2020) first introduce QMIX, but only consider tasks involving a limited number of drivers in simulated grid scenarios, which presents a significant gap to real-world applications. Hao et al. (Hao and Varakantham 2022) explore hierarchical structures to address the curse of dimensionality in QMIX (Rashid et al. 2020), but the input dimensions of the mixture network still increase with the number of agents. Bose et al. (Bose et al. 2023) propose a Value Decomposition (VD) (Sunehag et al. 2018) based method; however, the credit assignment challenge becomes more complex in ride-sharing scenarios with hundreds of agents.
@@
Lastly, there exists a series of works aimed at improving system efficiency by jointly considering relocation (Zhang et al. 2024), order bundling (Jiang, Xu, and Varakantham 2025), and multi-modal transportation (Hu, Dong, and Li 2025). Since these topics extend beyond the scope of this paper, we do not discuss them further here.
@@
Nevertheless, most methods rely on precise value-function estimation (e.g., Q/V-networks), which introduces bias and instability due to approximation errors (Hu, Feng, and Li 2025). Our work addresses this gap by proposing GRPO/OSPO—eliminating value estimation entirely through group rewards—for stable and scalable coordination in homogeneous MARL systems.

### v2

@@ 5 Conclusion
In this paper, we propose two novel MARL methods for the order dispatch task in the ride-sharing system with AVs that do not rely on estimating V-values or Q-values. First, we successfully adapt GRPO to the ride-sharing context, achieving promising performance by eliminating the estimation errors associated with the V-value network in PPO. Building on the insights gained from GRPO, we further introduce the OSPO method, which enables policy optimization using only single-step rewards. To the best of our knowledge, OSPO is the simplest and most efficient RL method for order dispatch on ride-sharing platforms to date. Extensive experiments on real-world ride-sharing data demonstrate that our methods consistently outperform previous approaches in most scenarios, while requiring only a simple MLP network and exhibiting low GPU utilization.
Impact Statement
We believe this work raises minimal ethical concerns, as the primary focus is on automated vehicles. However, in the future, we would like to explore whether our approach could benefit human drivers in ride-sharing systems. The core principle of our method is the homogeneous V-value (long-term profit), which may contribute to improving fairness in these systems.
Appendix A Algorithm
@@
The detailed algorithm of our GRPO/OSPO for ride-sharing system is illustrated in Algorithm 1.
Algorithm 1 GRPO/OSPO for Ride-Sharing System
@@
1:for episode =1=1 to $M$ do
2:  Initialize experience buffer $\mathcal{D}\leftarrow\emptyset$
@@
10:   Store transition $(\{s_{i,t}\},s^{o}_{t},U_{t},\{r_{i,t+1}\},\{s_{i,t+1}\},s^{o}_{t+1})$ in $\mathcal{D}$
@@
12:  for epoch =1=1 to $H$ do
13:   Sample batch $\mathcal{B}\sim\mathcal{D}$
14:   Compute group advantage $\mathcal{A}_{t}$
15:   Update policy parameters $\theta$ by maximizing Eq. 3
16:  end for
17:  if current policy $\pi_{\theta}$ outperforms $\pi_{\phi}$ then
18:   Update best checkpoint: $\phi\leftarrow\theta$
19:  end if
20:end for
Appendix B Experiment Details
B.1 Evaluation Metrics
The evaluation metrics used in this paper include:
  • Served Rate: The proportion of confirmed trips relative to the total number of trip requests initiated by customers.
  • Delivery Time: The total time taken to serve a trip from origin to destination.
  • Detour Time: The additional time spent on delivery beyond the minimum delivery time, which is defined as the time required if the vehicle serves only this trip without bundling others.
  • Pickup Time: The waiting time for customers between trip confirmation and the arrival of the vehicle at the trip origin.
  • Confirmation Time: The waiting time for customers from the moment they initiate a trip request to when the platform assigns the trip to a vehicle.
B.2 Detailed Experiment Results
Figure 4: Order Temporal Distribution
Figure 5: Experiment Results in Testing Set
In Fig. 4 and 5, we illustrated the order amount and the metrics of different methods under each hour in the whole day. First, it’s intuitive that the trend of reward is similar to the order volume, since more orders can bring more profits to platform. Then we notice before 8 a.m., when the order volume is very low, the overall performance of different methods is similar, since the idle vehicles are sufficient. However, after that, the BMG-Q and HIVES first illustrate slightly advantage during 8 a.m. to 1 p.m. and then are outperformed by our GRPO and OSPO methods as the order volume further increases after 1 p.m.. Overall, our GSPO and OSPO mainly outperform others by optimizing the pickup time and confirmation time, leading to the highest order service rate in the on-peak hour 6 p.m..
Additionally, to further prove our homogeneity assumption, we illustrate the relationship among initial state (location) and metrics regarding cumulative reward (without discount factor $\gamma$) and reward-to-go (with discount factor $\gamma$) of our method. The results suggest that in our method, the state does not significantly influence the long-term cumulative reward and reward-to-go. As shown in Fig. 6, this supports our homogeneity assumption.
Figure 6: The relationship between initial location and metrics about cumulative reward and reward-to-go.
B.3 Expanded Experiment
Table 2: Performance of Different Methods in Queens, New York City (TLCData)
Method Reward Service Rate Delivery Time Detour Time Pickup Time Confirmation Time DeepPool 5222.85 0.64 11.24 1.84 12.30 0.21 BMG-Q 5362.00 0.66 9.63 1.08 12.98 0.27 HIVES 3560.80 0.60 8.30 0.36 14.67 0.41 Enders et al. 4543.68 0.61 10.41 0.85 13.39 2.25 CEVD 4388.83 0.62 11.61 1.33 13.74 0.29 Hoppe et al. 4728.93 0.63 9.78 0.53 13.45 2.18 GRPO 5365.89 0.64 10.56 1.86 12.21 0.12 OSPO 5521.74 0.67 11.37 2.21 11.88 0.21
To further demonstrate the generalization of our method, we conduct experiments using High Volume For-Hire Vehicle (FHV) trip data from Queens, New York City (TLCData). Unlike the Manhattan data used in our previous experiment, Queens is a larger area with a lower order density. This difference in order distribution can complicate the process of determining which orders to bundle together for vehicle sharing. In this scenario, we select the data from 19:00 to 19:30 on July 17, 2024, which includes 2,024 valid orders and 500 vehicles. The detailed experimental results are presented in Table 2.
The results demonstrate that our proposed GRPO and OSPO methods continue to exhibit superior performance, primarily by optimizing pickup and confirmation times. Specifically, OSPO significantly outperforms the other methods, including GRPO. This advantage arises because, in a more challenging and dynamic scenario (due to the enlarged area), the GRPO method is more severely impacted by errors associated with the bounded time horizon, whereas OSPO leverages a one-step reward to mitigate this shortcoming.
Appendix C Related Work
@@
In this section, we provide a brief overview of current ride-sharing methods. Starting from the Independent Double Deep Q-Learning (IDDQN)-based DeepPool (al2019deeppool), DTDE methods have become a mainstream paradigm, as the large number of agents, along with the extensive state and action spaces, severely hinders the development of centralized methods. Subsequent works have primarily aimed to enhance algorithms by fostering better cooperation among agents. For example, Graph Neural Networks (GNNs) have been widely explored to effectively utilize neighboring information (hu2025bmg). li2019efficient also propose using mean field approaches to capture more useful global information.
@@
Additionally, some works attempt to transfer the successes of CTDE and CTCE to ride-sharing, where centralized critics can facilitate better cooperation among agents. However, these approaches face various shortcomings and challenges. de2020efficient first introduce QMIX to ride-sharing, but only consider tasks involving a limited number of vehicles in simulated grid scenarios, which presents a significant gap to real-world applications. hao2022hierarchical explore hierarchical structures to address the curse of dimensionality in QMIX (rashid2020monotonic), but the input dimensions of the mixture network still increase with the number of agents. bose2023sustainable propose a Value Decomposition (VD) (sunehag2018value) based method; however, the credit assignment challenge becomes more complex in ride-sharing scenarios with hundreds of agents.
@@
Lastly, there exists a series of works aimed at improving system efficiency by jointly considering relocation (zhang2024joint), order bundling (jiang2025optimizing), and multi-modal transportation (hu2025coordinating). Since these topics extend beyond the scope of this paper, we do not discuss them further here.
@@
Nevertheless, most methods rely on precise value-function estimation (e.g., Q/V-networks), which introduces bias and instability due to approximation errors (hu2025bmg). Our work addresses this gap by proposing GRPO/OSPO—eliminating value estimation entirely through group rewards—for stable and scalable coordination in homogeneous MARL systems.